# Context-Aware Coupler Reconfiguration for Tunable Coupler-Based Superconducting Quantum Computers

### v1

@@ 4.3 Fidelity Improvement
CA-CORE outperforms existing topologies in all evaluations. Particularly noteworthy is the 40.13% improvement in fidelity at an error rate ($\epsilon$) of 0.001 compared with the result achieved by Sycamore, as shown in Figure 6(b). On average for all error rate, CA-CORE exhibits fidelity improvements of 15.76%, 16.39%, 7.52%, 5.69%, 13.91%, and 9.77%, as shown in Figure 6(a),  6(b),  6(c),  6(d),  6(e), and 6(f), respectively.
Notably, in the result shown in Figure 6(d) for the ‘bv_n19’ circuit, the improvement is not significant at lower error rates. The ‘bv_n19’ circuit represents the Bernstein-Vazirani algorithm [38]. This algorithm requires a brief quantum circuit with two H gates for all qubits, one NOT (X) gate, and $n-1$ CNOT gates, where $n$ is the qubit count. Each CNOT gate connects all qubits with the last one. This structure ensures that at lower error rates, errors across different setups have minimal impact on result fidelity. Further, notably, Sycamore (24) performs worse compared to all the other topologies. This is primarily because the experiment was performed by employing basic qubit-mapping method. For the compiled circuit on various topologies, the circuit depths are as follows: 40 for CA-CORE, 43 for ibmq_almaden, 49 for ibmq_johannesburg, 48 for ibmq_cairo, 53 for Google’s Sycamore (24), and 45 for ibmq_prague. With differences in circuit depth, the impact of relatively high error rates also increases, contributing to the inferior performance of Google Sycamore’s topology. This contributes to the marginal improvement seen in CA-CORE compared to existing topologies across low-to-higher error-rate configurations.
@@ 4.4 Scalable Time Complexity
(a) ’multiplier_n15’
(b) ’Random Circuit 30-Qubit’
@@ 5 Discussion and Future Work
The manuscript refrains from directly comparing the proposed qubit mapping methodology with existing approaches due to a fundamental divergence in their focal points. Existing qubit mapping strategies predominantly concentrate on fixed topology target quantum hardware, aiming to optimize within the constraints of predetermined qubit layouts. In contrast, the manuscript’s emphasis lies in configuring optimal qubit coupling maps, specifically targeting tunable coupler-based quantum hardware. As a result, the comparison isn’t directly pursued as the existing qubit mapping techniques are tailored to a different paradigm, focusing on fixed hardware configurations, whereas the manuscript’s methodologies are designed for systems with tunable coupler features, warranting a distinct evaluation approach
@@ 6 Related Work
These studies primarily emphasize the design of tunable couplers to enhance the scalability of SQC processors. Gushu et al. [23] emphasized a logical qubit coupling topology and construct the hardware architecture using three subroutines: layout design, bus selection, and frequency allocation. This was achieved while considering physical constraints based on the profiling information from the coupling degree list and the coupling strength matrix. Their primary objective was to enhance the yield rate by maximizing qubit connectivity. However, they did not address the noise implications of their architecture. Tan et al. [43] presented QLSA-reconfigurable-atomic-arrays (RAA), a layout synthesizer RAA during the QC compilation phase. While their method achieves gate reduction and heightened fidelity for reconfigurable architectures through hardware-adaptive constraints, it is specifically designed for atomic arrays. Conversely, our proposed approach addresses the constraints inherent in SQCs.
Lin et al. [24] proposed a method similar to ours which utilizes layout synthesis to establish qubit edge connections. However, the edges they determined contradict the fabrication constraints due to the frequency collision outlined in [22, 23]. Conversely, our proposed method considers the edge connection that adheres to the frequency collision constraint, maximizing the SQC’s connectivity. Given these contradictions and disparities in architectural design for edge selection, we cannot conduct a comparative study between their method and our method.

### v2

@@ 4.2 Fidelity Improvement
CA-CORE outperforms existing topologies in all evaluations. Particularly noteworthy is the 40.13% improvement in fidelity at an error rate ($\epsilon$) of 0.001 compared with the result achieved by Sycamore, as shown in Figure 7(b). On average for all error rates, CA-CORE exhibits fidelity improvements of 15.76%, 16.39%, 7.52%, 5.69%, 13.91%, and 9.77%, as shown in Figure 7(a),  7(b),  7(c),  7(d),  7(e), and 7(f), respectively.
Notably, in the result shown in Figure 7(d) for the ‘bv_n19’ circuit, the improvement is not significant at lower error rates. The ‘bv_n19’ circuit represents the Bernstein-Vazirani algorithm [38]. This algorithm requires a brief quantum circuit with two H gates for all qubits, one NOT (X) gate, and $n-1$ CNOT gates, where $n$ is the qubit count. Each CNOT gate connects all qubits with the last one. This structure ensures that at lower error rates, errors across different setups have minimal impact on result fidelity. Further, Google Sycamore (24) performs worse than all the other topologies. This is primarily because the experiment employed a trivial qubit-mapping method. For the compiled circuit on various topologies, the circuit depths are as follows: 40 for CA-CORE, 43 for ibmq_almaden, 49 for ibmq_johannesburg, 48 for ibmq_cairo, 53 for Google’s Sycamore (24), and 45 for ibmq_prague. With differences in circuit depth, the impact of relatively high error rates also increases, contributing to the inferior performance of Google Sycamore’s topology. This contributes to the marginal improvement seen in CA-CORE compared to existing topologies across low-to-higher error-rate configurations.
@@ 4.3 Scalable Time Complexity
In the reprocessing step, an input graph of $n$ qubits is prioritized, where each qubit can connect to any other. This phase identifies all qubit correlations and their coupling weights with a time complexity of $O(n^{2})$. Then, Algorithm 1 constructs a new graph to maximize the weighted path using the obtained qubit connections from the preprocessing phase.
@@
(a) ‘multiplier_n15’
(b) ‘Random Circuit 30-Qubit’
@@ 5 Discussion and Future Work
The manuscript refrains from directly comparing the proposed qubit mapping methodology with existing approaches due to a fundamental divergence in their focal points. Existing qubit mapping strategies predominantly concentrate on fixed topology target quantum hardware, aiming to optimize within the constraints of predetermined qubit layouts. In contrast, the manuscript emphasizes configuring optimal qubit coupling maps, specifically targeting tunable coupler-based quantum hardware. As a result, the comparison isn’t directly pursued as the existing qubit mapping techniques are tailored to a different paradigm, focusing on fixed hardware configurations, whereas the manuscript’s methodologies are designed for systems with tunable coupler features, warranting a distinct evaluation approach
@@ 6 Related Work
These studies primarily emphasize the design of tunable couplers to enhance the scalability of SQC processors. Gushu et al. [23] emphasized a logical qubit coupling topology and constructed the hardware architecture using three subroutines: layout design, bus selection, and frequency allocation. This was achieved while considering physical constraints based on the profiling information from the coupling degree list and the coupling strength matrix. Their primary objective was to enhance the yield rate by maximizing qubit connectivity. However, they did not address the noise implications of their architecture. Tan et al. [43] presented QLSA-reconfigurable-atomic-arrays (RAA), a layout synthesizer RAA during the QC compilation phase. While their method achieves gate reduction and heightened fidelity for reconfigurable architectures through hardware-adaptive constraints, it is specifically designed for atomic arrays. Conversely, our proposed approach addresses the constraints inherent in SQCs.
Lin et al. [24] proposed a method similar to ours which utilizes layout synthesis to establish qubit edge connections. However, the edges that they determined contradict the fabrication constraints due to the frequency collision outlined in [22, 23]. Conversely, our proposed method considers the edge connection that adheres to the frequency collision constraint, maximizing the SQC’s connectivity. Given these contradictions and disparities in architectural design for edge selection, we cannot conduct a comparative study between their method and our method.